OpenStudy (dj3strella):

Question 2.2. Select the proper equation in point-slope form for the given point and slope. Point: (–8, 3); Slope: 6

OpenStudy (anonymous):

Point slope form! y – y1 = m(x – x1) You have a y coordinate to plug in for y1, an x coordinate to plug in for x1 and a slope value to plug in for m. Can you take it from there?

OpenStudy (dj3strella):

y + 3 = 6x – 48 y + 3 = 6(x + 8) y – 3 = 6(x – 8) * y – 3 = 6(x + 8) is the answer *